a) Eddie Corp has a cm/unit = $13.80 and a tax rate = 25%. ATNI currently is $53,820. What is ATNI if Eddie sells an additional 800 units?

b) Roger has a DM efficiency variance of $22,050 unfavorable. Roger actually used 2.2 pounds/unit to produce 42,000 units. Assume the standard cost per pound = $1.75. What are standard pounds per unit?
Answer #1
Eddie Corp
New ATNI = Current ATNI + ( Additional units * cm/unit ) * ( 1 - Tax% ) = 53820 + ( 800 * 13.80 ) * ( 1 - 25% ) 62100
Roger
Actual quantity = Actual quantity/unit * Production = 2.2 * 42000 92400
DM efficiency variance = ( Actual quantity - Standard quantity ) * Standard rate
22050 = ( 92400 - Standard quantity ) * 1.75
92400 - Standard quantity = 22050 * 1.75
92400 - Standard quantity = 38587.50
Standard quantity = 92400 - 38587.50
Standard quantity = 53812.5
Standard pounds per unit = Standard quantity / Production = 53812.5 / 42000 1.28 pounds/unit